OpenStudy (anonymous):

find the intercepts of f(x)=3x^2+2x-5

OpenStudy (anonymous):

(x - 1 ) (3x +5) = 0 x = 1 x = -5/3

OpenStudy (anonymous):

i got this one can you help me with \[f(x)=\sin ^2x+\cos ^2x, x=9\pi/4\]

OpenStudy (anonymous):

sin^2 (x) + cos^2 (x) = 1, always, so f(x) = 1. f(9pi/4) = (sin 9pi/4)^2 + (cos 9pi/4)^2 = (sin 45°)^2 + (cos45°)^2 = (sqrt2/2)^2 + (sqrt2/2)^2 = 2/4 + 2/4 = 4/4 = 1